Use the graph to determine a. the​ function's domain; b. the​ function's range; c. the​ x-intercepts, if​ any; d. the​ y-intercept, if​ any; and e. the missing function​ values, indicated by question​ marks, below. f(-2)=? f(2)?

A coordinate system has a horizontal x-axis labeled from negative 5 to 5 in increments of 1 and a vertical y-axis labeled from negative 5 to 5 in increments of 1. A parabola that opens upward has vertex (negative 2, negative 1) and passes through the points (negative 4, 0), (0, 0), and left parenthesis 2 comma 3 right parenthesis.

a. The domain is
b. The range is
c. Select the correct choice below and fill in any answer boxes within your choice.
A.The​ x-intercept(s) is​ (are)
d. Select the correct choice below and fill in any answer boxes within your choice.
A.The​ y-intercept is
f(-2)=
f(2)=
1
Expert's answer
2020-09-02T17:58:05-0400

a. The curve is parabola with the formula y=ax2+bx+c , so x can be any number and graph domain is "(-\infty, +\infty)" .

b. As you can see from the graph, y changes from "+\infty" to -1 in the point (-2,-1) and then again to +"\infty". So range is "[-1, +\infty)."

c. Points B(-4, 0) and C(0, 0) are points with x-intercepts. In these points y=0.

d. Point C(0, 0) is y-intercept. In this point x=0.

f(-2)=-1 because of coordinates of point A

f(2)=3 because of coordinates of point D